construct an ellipse when the distance of its focus from its directrix is equal to 60 mm and the eccentricity is 2/3.also draw a tangent and a normal to the curve
Answer by stanbon(75887)

what is 7 2/5 divide by 2 7/9?
------
7 2/5 / 2 7/9
------
= (5*7+2)/5 / (9*2+7)/9
------
= 37/5 / 25/9
----
= (37*9)/(5*25)
---
= 363/125
----
Cheers,
Stan H.
===========